OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. A.y – 2 = 3(x – 5) B.y + 8 = 3(x + 5) C.y – 5 = 3(x – 2) D.y + 5 = 3(x + 2)

OpenStudy (anonymous):

C. Point slope form is: \[y - y _{1} = (slope)(x-x_{1})\]

OpenStudy (anonymous):

y + 5 = 3(x + 2) im guessing its d?